OpenStudy (anonymous):

find the real roots of the equaton 18/x^4+1/x^2=4

OpenStudy (anonymous):

The Real roots of the given equation is X = 3/2 and -3/2 Multiply with x^4 in the whole equation Then there will be a equation = 4x^4 - x^2 - 18 = 0 On solving, u will get equation = ( x^2 + 2)(4x^2 - 9) = 0 Solve for the real part & and u will get the Answer..!!

OpenStudy (anonymous):

how did u get ( x^2 + 2)(4x^2 - 9

OpenStudy (anonymous):

@Trexy

OpenStudy (anonymous):

wait.. XD

OpenStudy (anonymous):

Use ac test to factor. First, you factor 4 into 4 and 1, -18 into -9 and 2 as below 4 -9 1 2 Then you cross multiply to make sure you have -1. Then add x to the right column to write (4x^2 - 9)(x^2 +2)

OpenStudy (anonymous):

Sorry about that.. anyway.. he got the answer...

OpenStudy (anonymous):

how do u factor out 4 into four into1?um a bit puzzled

OpenStudy (anonymous):

Hahah.. ok wait.

OpenStudy (anonymous):

@Trexy

OpenStudy (anonymous):

@Mr._To

OpenStudy (anonymous):

because 4*1=4 which gives your x^4 coefficient. The 2 and -9 give you back your -18 constant. Now all you need is the coefficient for x^2. you're trying to determine if the 4 and 1, -9 and 2, factorization will give you back your -1 (for the x^2) upon expansion. It does. So you end up with the 4 and 1 attached to your x^2 terms and 2 and -9. (4x^2-9)(x^2+2)=0

OpenStudy (anonymous):

@Tushara excellent solution!

OpenStudy (anonymous):

:) i hope that helped

OpenStudy (anonymous):

Thanks @Tushara and @eseidl helped a lot :)

OpenStudy (anonymous):

m glad, np